# Sub-diffraction estimation, discrimination and learning of quantum states of light

### v1

@@ III The special role played by the HG basis
From this we obtain the density matrix in Eq. (4), conditioned on having at least one photon (i.e., we are neglecting the vacuum contribution),
$\rho$ $=\sum_{x,y}I_{x,y}|T_{x,y}\rangle\langle T_{x,y}|$ (17)
$=\sum_{x,y}I_{x,y}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}\sum_{m,n}\frac{1}{\sqrt{m!n!}}\frac{x^{m}y^{n}}{(2\sigma)^{m+n}}|\text{HG}_{mn}\rangle$ (18)
×$\phantom{=}~{}\times\sum_{m^{\prime},n^{\prime}}\frac{1}{\sqrt{m^{\prime}!n^{\prime}!}}\frac{x^{m^{\prime}}y^{n^{\prime}}}{(2\sigma)^{m^{\prime}+n^{\prime}}}\langle\text{HG}_{m^{\prime}n^{\prime}}|$ (19)
@@
Consider now the sub-Rayleigh regime. In this limit we can truncate the expansion up to the second order in $x/\sigma$ and $y/\sigma$. This yields
$\rho$ $\simeq\left(1-\frac{\langle x^{2}\rangle+\langle y^{2}\rangle}{4\sigma^{2}}\right)|\text{HG}_{00}\rangle\langle\text{HG}_{00}|$
$\phantom{=}~{}+\frac{\langle x\rangle}{2\sigma}\left(|\text{HG}_{00}\rangle\langle\text{HG}_{10}|+|\text{HG}_{10}\rangle\langle\text{HG}_{00}|\right)$
$\phantom{=}~{}+\frac{\langle y\rangle}{2\sigma}\left(|\text{HG}_{00}\rangle\langle\text{HG}_{01}|+|\text{HG}_{01}\rangle\langle\text{HG}_{00}|\right)$
$\phantom{=}~{}+\frac{\langle x^{2}\rangle}{4\sigma^{2}}|\text{HG}_{10}\rangle\langle\text{HG}_{10}|+\frac{\langle y^{2}\rangle}{4\sigma^{2}}|\text{HG}_{01}\rangle\langle\text{HG}_{01}|$
$\phantom{=}~{}+\frac{\langle xy\rangle}{4\sigma^{2}}\left(|\text{HG}_{10}\rangle\langle\text{HG}_{01}|+|\text{HG}_{01}\rangle\langle\text{HG}_{10}|\right)\,,$ (21)
where $\langle x^{2}\rangle:=\sum_{x,y}I_{x,y}x^{2}$, $\langle xy\rangle:=\sum_{x,y}I_{x,y}xy$, etc.
This expansion shows that in the sub-Rayleigh regime the state is completely determined by the first and second moments of the intensity distribution.
A further simplification is obtained if we assume that the optical system is aligned towards the “center of mass” (defined using the intensity as weights) of the source object. This means that the first moments vanish,
$\langle x\rangle=\langle y\rangle=0\,,$ (22)
and the state in Eq. (21) becomes block-diagonal, that is, the cross-terms $|\text{HG}_{00}\rangle\langle\text{HG}_{10}|$, $|\text{HG}_{00}\rangle\langle\text{HG}_{01}|$ are suppressed.
In conclusion, the state of a single photon emitted by the source, in the sub-Rayleigh regime and if the optical system is aligned to the “center of mass”, is completely determined by the covariance matrix of the intensity distribution,
$\gamma=\sum_{x,y}I_{x,y}\left(\begin{array}[]{cc}x^{2}&xy\\ xy&y^{2}\end{array}\right)\,.$ (25)
Note that photon detection in the modes $\text{HG}_{01}$ and $\text{HG}_{10}$ directly estimates the second order moments. To estimate the off-diagonal terms we need to rotate the HG modes. A rotation of the optical system of an angle $\varphi$ defines the rotated HG modes
$|\text{HG}_{01}(\varphi)\rangle$ $=\cos{\varphi}|\text{HG}_{01}\rangle-\sin{\varphi}|\text{HG}_{10}\rangle$ (26)
$|\text{HG}_{10}(\varphi)\rangle$ $=\sin{\varphi}|\text{HG}_{01}\rangle+\cos{\varphi}|\text{HG}_{10}\rangle\,.$ (27)
For a particular choice of $\varphi$ the covariance matrix becomes diagonal. Therefore, photon detection in the lower HG modes, rotated by a suitable angle, completely determines the source in the sub-Rayleigh regime.
III.1 Symmetric binary discrimination
In binary classification, two source objects are given, which determine two corresponding states of light, $\rho_{0}$ and $\rho_{1}$. As we have seen above, in the sub-Rayleigh regime these states are equivalently represented by the intensity covariance matrices $\gamma_{0}$ and $\gamma_{1}$.
In the framework of symmetric discrimination, one considers the average probability of obtaining a false positive or a false negative. The fundamental limit on such a probability is given by the Helstrom bound Helstrom (1976), which expresses the minimum average error probability in terms of the trace distance between the states $\rho_{0}$ and $\rho_{1}$, or equivalently between the covariance matrices $\gamma_{0}$ and $\gamma_{1}$,
$P_{\text{err}}=\frac{1-\Delta}{2}\,,$ (28)
where
$\Delta=\frac{1}{2}\|\rho_{0}-\rho_{1}\|_{1}=\frac{\varepsilon}{8\sigma^{2}}\left(\delta+\|\gamma_{0}-\gamma_{1}\|_{1}\right)\,,$ (29)
and $\|\,\cdot\,\|_{1}=\mathrm{Tr}|\,\cdot\,|$ denotes the $1$-norm, and
$\delta=\left|\mathrm{Tr}\gamma_{0}-\mathrm{Tr}\gamma_{1}\right|\,.$ (30)
It is known that the optimal measurement that saturates the Helstrom bounds corresponds to projections into the positive and negative parts of the difference $\rho_{0}-\rho_{1}$, which in our setting is equivalent to the projectors onto the positive and negative parts of $\gamma_{0}-\gamma_{1}$. Since $\gamma_{0}-\gamma_{1}$ is a $2\times 2$ matrix, each of these projector is a one-dimensional projector corresponding to the eigenvectors of $\gamma_{0}-\gamma_{1}$. We can then distinguish two cases:
  1. 1.
    If the matrices $\gamma_{0}$ and $\gamma_{1}$ commute, then they share a common pair of eigenvectors, which are obtained for a particular choice of the rotation angle $\varphi$ in Eqs. (26)-(27). Such value of the angle is determined by the common axes of symmetry of the objects. In turn, the projectors into the positive and negative part of $\gamma_{0}-\gamma_{1}$ corresponds to rotated HG states and the optimal measurement is HG detection along the angle $\varphi$.
  2. 2.
    If the matrices do not commute, we need to consider the eigenvectors of $\gamma_{0}-\gamma_{1}$. Without loss of generality, the eigenvectors are once again of the form in Eqs. (26)-(27) for some specific value of $\varphi$. Therefore also in this case detection in rotated HG modes is optimal.
In conclusion, we have obtained the following result concerning sub-Rayleigh image discrimination:
Proposition 1
Photon detection in the rotated HG modes saturates the Helstrom bound for ultra-weak signals in the sub-Rayleigh regime.
III.2 Estimation of the moments
Consider a quantum state $\rho(\vartheta)$ depending on an unknown parameter $\vartheta$. The goal of parameter estimation is to estimate $\vartheta$ by measuring $N$ instances of the state.
In the context of imaging, Tsang and collaborators showed in Ref. Tsang et al. (2016) that DI is not suitable for parameter estimation when the size of the object is much smaller than the Rayleigh length. They showed that the statistical error in the estimation becomes unbounded in the sub-Rayleigh limit. This phenomenon, dubbed the Rayleigh curse, is the manifestation of the Rayleigh resolution criterion in the context of parameter estimation.
In the same work, Tsang and collaborators also showed that HG SPADE lifts the Rayleigh curse, allowing parameter estimation with a statistical error that does not depend on the size of the object. HG SPADE also saturates the ultimate quantum limit as expressed by the quantum Cramér-Rao bound. The problem considered in Ref. Tsang et al. (2016) was the estimation of the transverse separation $s$ between two point-like sources. Similar conclusions were also drawn by Zhou and Jiang Zhou and Jiang (2019) for the estimating of the second moments of the intensity distribution of a two-dimensional incoherent source laying in the object plane.
Consider the moments of order $k$:
$M_{i,j}=\sum_{x,y}I_{x,y}x^{i}y^{j}\,,$ (31)
with $i+j=k$, and define the parameters
$\vartheta_{i,j}:=\sqrt{M_{i,j}}$ (32)
as the square root of the moments. Note that for the case of two point-sources of equal intensity, with transverse separation $s$ along the $x$-axis, and if the optical system is aligned to the mid-point, we obtain $s^{2}=4M_{2,0}$ and $s=2\vartheta_{2,0}$. Therefore, the estimation of the parameters $\vartheta_{i,j}$ includes the original problem of Tsang and collaborators as a special case.
We now show why HG SPADE allows us to lift the Rayleigh curse for the estimation of the separation $s$, and more generally for the estimation of the parameters $\vartheta_{i,j}$. Consider first the estimation of $\vartheta_{2,0}=\sqrt{\langle x^{2}\rangle}$. According to the decomposition in HG basis in Eq. (21), the moments of order two are proportional to the probability of detecting a photon in the lower HG modes. For example, photon detection in $\text{HG}_{10}$ has probability
$p_{10}=\frac{\varepsilon}{4\sigma^{2}}\langle x^{2}\rangle\,.$ (33)
It follows from the properties of the binomial distribution that, given $N$ photon detection events, the statistical error in the estimation of $p_{10}$ is
$\Delta p_{10}=\sqrt{\frac{p_{10}(1-p_{10})}{N}}\,,$ (34)
which also saturates the Cramér-Rao bound. In the sub-Rayleigh limit, we have $p_{10}\ll 1$, therefore
$\Delta p_{10}\simeq\sqrt{\frac{p_{10}}{N}}\,.$ (35)
By error propagation, from this we obtain the statistical error in the estimation of $\sqrt{p_{10}}$:
$\Delta\sqrt{p_{10}}\simeq\frac{1}{2}\sqrt{\frac{1}{N}}\,,$ (36)
and finally the error in the estimation of the parameter $\vartheta_{2,0}=2\sigma\sqrt{p_{10}/\varepsilon}$:
$\Delta\vartheta_{2,0}\simeq\sigma\sqrt{\frac{1}{\varepsilon N}}\,.$ (37)
(Note that $\varepsilon N$ is the average number of photons detected.)
We can actually generalise this result. From the HG expansion in Eq. (20) we see that photon detection in modes $\text{HG}_{mn}$ is related to the even moments $\langle x^{2m}y^{2n}\rangle$. Reasoning as above we conclude that the Rayleigh curse is lifted for the estimation of the parameters $\vartheta_{2m,2n}=\sqrt{\langle x^{2m}y^{2n}\rangle}$.
In conclusion, we have obtained the following result:
Proposition 2
Photon detection in HG modes lifts the Rayleigh curse for the estimation of the square root of all even moments, $\vartheta_{2m,2n}=\sqrt{\langle x^{2m}y^{2n}\rangle}$. Furthermore, the parameters $\vartheta_{2m,2n}$ can be jointly estimated with a single HG SPADE measurement.
Figure 3: Statistical error $\Delta\sqrt{\langle x^{2}\rangle}$ obtained from simulated $N=10^{3}$ photon detection events, plotted versus the noise parameter $\tilde{\sigma}=\sigma/f$, where $\sigma=9.5$ pixels and $f$ is a scaling factor. The blue dots are for DI, where the square root of the second moment $\vartheta_{2,0}=\sqrt{\langle x^{2}\rangle}$ is computed using Eq. (38). The orange squares are for SPADE, where the estimated quantity is $\sqrt{p_{10}}$. We note that the error from DI increases with increasing $\tilde{\sigma}$, whereas the error from SPADE remains constant.
Since it allows for a direct estimation of (the square root of) the even moments, it is natural to expect HG SPADE to outperform DI for this task. In fact, to obtain the moments from DI we first need photon detection in each pixel to estimate the quantities $p_{xy}^{\text{DI}}$ in Eq. (7). For example, to estimate the quantity $\langle x^{2}\rangle$, we need to combine them in the formula:
$\langle x^{2}\rangle=\sum_{x,y}x^{2}p_{xy}^{\text{DI}}-\sigma^{2}\,,$ (38)
where the terms $\sigma^{2}$ account for de-convolution of the Gaussian blurring. As errors propagate and accumulate, we expect a lower signal-to-noise ratio when estimating $\vartheta_{2,0}$ using data collected by DI.
We have numerically verified this by considering an object from the MNIST data set (the one in Fig. 2), re-scaling it by a factor $1/f$, and then applying a Gaussian noise with width $\sigma$. The effects of diffraction are quantified by quantity $\tilde{\sigma}=\sigma/f$. Since the original MNIST image is pixelated, aliasing errors occur when the image is re-scaled. We have selected a range of values of $f$ such that the aliasing error has only a limited impact. A comparison between the statistical error $\Delta\sqrt{\langle x^{2}\rangle}$ in the estimation of $\vartheta_{2,0}=\sqrt{\langle x^{2}\rangle}$ with DI and SPADE is shown in Fig. 3.
@@ IV Numerical experiments
In this Section we present a number of numerical experiments where HG SPADE is combined with machine learning algorithms into a hybrid quantum-classical approach to classify faint sources in the sub-Rayleigh regime. The dataset used for the experiments is the MNIST, handwritten digits, which is extracted from the torchvision Python library. This is a widesrpread benchmark dataset in computer vision and machine learning.
First we simulate the action of the optical imaging system, which focuses the field emitted from the object sources. For experiments of binary classification, we consider restricted datasets of either 00’s and $1$’s (in Section IV.1) or $6$’s and $9$’s (in Section IV.1). All classes of MNIST objects, i.e. the digits from zero to nine, are considered in Section IV.2. Diffraction through the pupil of the optical system induces blurring. In our simulation we fix the value of the width of the PSF to $\sigma=9.5$ (measured in number of pixels), and scale the source object by a factor $f$: this induces a relative width of $\tilde{\sigma}=\sigma/f$. From the focused optical field we can compute the probability of photon detection for both DI and HG SPADE, respectively using Eq. (7) and Eq. (12). The statistics for given number of detected photon $N$ is generated by Monte Carlo simulation.
Finally, the datasets obtained in this way are directly ingested into a machine learning model trained for classification. The models used in the following for performing the classification are standard ones in the machine learning literature: Random Forest Breiman (2001) and Fully Connected Neural Network (FCNN) Goodfellow et al. (2016). RF is an ensemble learning methods that proves to be quite effective both in regression and in classification. Furthermore, it is relativity light, hence, given the size of the HG SPADE vector, it is expected to capture the fundamental features useful for the classification. FCNN on the contrary is a deep learning method, particularly used for classification, that extracts features from the input sample via a series of connected hidden layers with non-linear activation functions. It is one of the standard models for automated classification, particularly useful in computer vision.
IV.1 Binary classification
@@
Note that the MNIST objects are already centered in such a way that $\langle x\rangle=\langle y\rangle=0$.
IV.1.1 SPADE
We have seen that SPADE allows us to optimally estimate the second moments. Therefore, we expect SPADE to be a good measurement choice for this particular classification problem.
We focus on SPADE in the diagonal basis, since moments in this basis appear to be more effective in separating the two classes, as shown in Fig. 4. The probability of detecting a photon in the lower diagonal modes $\text{HG}_{10}$, $\text{HG}_{01}$ can be computed directly (up to normalisation) from Eq. (20). Without taking the limit of $x/\sigma,y/\sigma\ll 1$ we obtain
$p_{00}^{\text{SPADE}}$ $=\sum_{x,y}I_{x,y}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}\,,$ (43)
$p_{10}^{\text{SPADE}}$ $=\frac{1}{8\sigma^{2}}\sum_{x,y}I_{x,y}(x+y)^{2}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}\,,$ (44)
$p_{01}^{\text{SPADE}}$ $=\frac{1}{8\sigma^{2}}\sum_{x,y}I_{x,y}(x-y)^{2}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}\,.$ (45)
@@ IV.1.1 SPADE
For any given integer value for the photon number $N$, we have implemented a MonteCarlo simulation to generate the relative frequencies $f_{uv}^{\text{SPADE}}(N)$ (see Section II.2) by sampling (with replacement) from the above distribution $p_{uv}^{\text{SPADE}}$. We have tested a RF using these relative frequencies as feature vectors. With a sample of $11867$ images from the database, and using $70\%$ of them for training and $30\%$ for validation. The results are displayed in Fig. 5, which shows the accuracy vs the PSF width $\tilde{\sigma}$, for different values of the photon number $N$.
@@
Figure 5 shows the qualitative behaviours of the accuracy: (1) the accuracy decreases with increasing $\tilde{\sigma}$, as the Gaussian noise induced by diffraction makes more difficult to discriminate the objects; (2) the accuracy increases with increasing $N$, due to smaller statistical fluctuations.
Figure 5: Accuracy for binary classification of 00’s and $1$’s using RF, performed on the HG SPADE vectors, evaluated in a rotated basis. The results are evaluated for different effective blurring factors $\tilde{\sigma}=9.5/f$ and considering various values of the number of photon detection events, $N$ (from bottom to topo, $N=100,500,1000,5000$). The error bars are calculated via stratified $k$-fold Ojala and Garriga (2010), using $m=10$ splits of the test dataset.
@@
Figure 6: Accuracy for the binary classification of 00’s and $1$’s, performed on the simulated DI (a) with RF, (b) with FCNN, (c) with RF on second moments computed from DI data. The results are evaluated for different effective blurring factors $\tilde{\sigma}=9.5/f$ and considering various photon detection events, $N$ (from bottom to top, $N=100,500,1000,5000$). The error bars are calculated via stratified $k$-fold Ojala and Garriga (2010), using $m=10$ splits of the test dataset.
IV.1.2 DI
@@
To compare with SPADE, we have performed a similar analysis for DI. For any number $N$ of photon detection events, we have generated the relative frequencies $f_{xy}^{\text{DI}}(N)$ (see Section II.1) by sampling (without replacement) from the DI probability distribution in Eq. (7). We have tested the machine learning approaches, with results displayed in Fig. 6. The top panel shows the results obtained for RF using the relative frequencies as feature vectors. The middles panel is for a FCNN, also using the relative frequencies as feature vectors. A direct comparison suggests that the FCNN performs much better than RF for this problem, as the hidden neurons are able to extract relevant features out of the original sample. We argue that, in particular, the FCNN is able to obtain the information most likely related to the moments of order two, which is crucial to classify correctly in the sub-Rayleigh regime.
@@
Finally, also for a fair comparison with SPADE, we helped the algorithm by performing a pre-computation of the second moments. Starting from the relative frequencies $f_{xy}^{\text{DI}}(N)$, the second moments can be estimated as follows:
$\langle x^{2}\rangle$ $\leftarrow\sum_{x,y}f_{xy}^{\text{DI}}(N)x^{2}\,,$ (46)
$\langle y^{2}\rangle$ $\leftarrow\sum_{x,y}f_{xy}^{\text{DI}}(N)y^{2}\,,$ (47)
@@
$\frac{1}{2}\langle(x+y)^{2}\rangle$ $\leftarrow\frac{1}{2}\sum_{x,y}f_{xy}^{\text{DI}}(N)(x+y)^{2}\,,$ (48)
$\frac{1}{2}\langle(x-y)^{2}\rangle$ $\leftarrow\frac{1}{2}\sum_{x,y}f_{xy}^{\text{DI}}(N)(x-y)^{2}\,,$ (49)
for the diagonal one. In particular, we have used the estimates of the second moments in diagonal basis and used them as feature vectors in a RF. The results are displayed in the bottom panel of Fig. 6, showing an increase in the accuracy compared to the FCNN fed with the relative frequencies. Moreover, it can be noticed that the fluctuations for each $\tilde{\sigma}$ are more contained compared to the previous cases: this is due to the fact that the second moments, which now are structured within the feature vector, are independent of $\tilde{\sigma}$ and hence they stabilise the computation.
In conclusion, DI supported by either RF and FCNN yields poor performances when compared against SPADE supported by RF. The accuracy of DI increases when the second moments are pre-computed from the DI relative frequencies and fed into a RF. This latter approach mimics the pre-processing implemented by SPADE, though in a physical rather than computational way. However, also this latter approach performs poorly in comparison with SPADE, especially for low photon number. This is due to the fact that computing the moments from the data obtained by DI yields a signal-to-noise ratio smaller than a direct measurement of SPADE Santamaria et al. (2023).
Figure 7: Confusion matrix of the output of the RF for multi-class classification, performed on the HG SPADE vectors. In particular the figure refers to the case with an effective blurring factor $\tilde{\sigma}=10$ and with number of photon detection events $N=5000$.

### v2

@@ III Expansion in the HG basis
From this we obtain the density matrix in Eq. (8) conditioned on having at least one photon (i.e., we are neglecting the vacuum contribution)
$\tilde{\rho}$ $=\sum_{x,y}I_{x,y}|T_{x,y}\rangle\langle T_{x,y}|$ (21)
@@ IV Numerical experiments
In this Section we present a number of numerical experiments where HG SPADE, or a suitable modification thereof, is combined with machine learning algorithms into a hybrid approach to classify faint sources in the sub-Rayleigh regime. A working code for the preprocessing hereby performed can be found at  [^2] https://github.com/PeppeGoodhelp/Spade_Estimation/. The dataset used for the experiments is the MNIST, handwritten digits, which is extracted from the torchvision Python library. This is a widespread benchmark dataset in computer vision and machine learning.
First we simulate the action of the optical imaging system, which focuses the field emitted from the object sources. For experiments of binary classification, we consider restricted datasets of either 00’s and $1$’s (in Section IV.1) or $6$’s and $9$’s (in Section IV.3). All classes of MNIST objects, i.e. the digits from zero to nine, are considered in Sections IV.2 and IV.3. Diffraction through the pupil of the optical system induces blurring. In our simulation we fix the value of the width of the PSF to $\sigma=9.5$ (measured in number of pixels), and scale the source object by a factor $f$: this induces a relative width of $\tilde{\sigma}=\sigma/f$. From the focused optical field we can compute the probability of photon detection for both DI and HG SPADE, respectively using Eq. (11) and Eq. (17). The statistics for given number of detected photon $N$ is generated by Monte Carlo simulation.
Finally, the datasets obtained in this way are directly ingested into a machine learning model trained for classification. The models used in the following for performing the classification are standard ones in the machine learning literature: Random Forest breiman2001 and Fully Connected Neural Network (FCNN) Goodfellow . RF is an ensemble learning method that proves to be quite effective both in regression and in classification. Furthermore, it is relativity light, hence, given the size of the HG SPADE vector, it is expected to capture the fundamental features useful for the classification. FCNN on the contrary is a deep learning method, particularly used for classification, that extracts features from the input sample via a series of connected hidden layers with non-linear activation functions. It is one of the standard models for automated classification, particularly useful in computer vision.
IV.1 Binary classification, of 00’s vs $1$’s
@@
Note that the MNIST objects are centered in such a way that we can put $\langle x\rangle=\langle y\rangle=0$. In our numerical experiment we focus on extracting information about the moments in the diagonal basis, as they appear to better separate the two classes.
IV.1.1 SPADE:
From Eq. (22) we see that information about the second moments in Eqs. (25)-(26) can be obtained through a measurement in the basis of rotated lower HG modes:
$\left\{|\text{HG}_{00}\rangle,\frac{|\text{HG}_{10}\rangle+|\text{HG}_{01}\rangle}{\sqrt{2}},\frac{|\text{HG}_{10}\rangle-|\text{HG}_{01}\rangle}{\sqrt{2}}\right\}\,.$ (27)
The probability of detecting a photon in these modes can be computed directly (up to normalisation) from Eq. (22):
$|\text{HG}_{00}\rangle$ $\to\sum_{x,y}I_{x,y}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}=p_{00}^{\text{SPADE}}\,,$ (28)
$\frac{|\text{HG}_{10}\rangle\pm|\text{HG}_{01}\rangle}{\sqrt{2}}$ $\to\frac{1}{8\sigma^{2}}\sum_{x,y}I_{x,y}(x\pm y)^{2}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}$
$\phantom{\to}=p_{\pm}^{\text{SPADE}}\,.$ (29)
@@ IV.1.1 SPADE:
For any given integer value for the photon number $N$, we have implemented a Monte Carlo simulation to generate the relative frequencies (see Section II.2) by sampling (with replacement) from the above probability mass distributions. We have tested a RF using these relative frequencies as feature vectors. With a sample of $11867$ images from the database, and using $70\%$ of them for training and $30\%$ for validation. The results are displayed in Fig. 4(d), which shows the accuracy vs the PSF width $\tilde{\sigma}$, for different values of the photon number $N$.
@@
Figure 4(d) shows the qualitative behaviours of the accuracy: (1) the accuracy decreases with increasing $\tilde{\sigma}$, as the Gaussian noise induced by diffraction makes more difficult to discriminate the objects; (2) the accuracy increases with increasing $N$, due to smaller statistical fluctuations.
@@
Figure 4: Accuracy in the binary classification of 00’s and $1$’s in the MNIST dataset. The four panels are obtained from different data and using different machine learning algorithms: (a) simulated DI data processed with RF; (b) simulated DI data processed with FCNN; (c) second moments computed from simulated DI data, processed with RF; (d) simulated SPADE data, evaluated in a rotated basis as in Eq. (29), processed with RF. The results are evaluated for different effective blurring factors $\tilde{\sigma}=9.5/f$ and considering various values of the number of photon detection events, $N$ (from bottom to topo, $N=100,500,1000,5000$). The error bars are calculated via stratified $k$-fold stratifiedk , using $m=10$ splits of the test dataset.
IV.1.2 DI:
@@
To compare with SPADE, we have performed a similar analysis for DI. For any number $N$ of photon detection events, we have generated the relative frequencies $f_{x^{\prime}y^{\prime}}^{\text{DI}}(N)$ (see Section II.1) by sampling (with replacement) from the DI probability distribution in Eq. (11). We have tested the machine learning approach, with results displayed in Fig. 4, panels (a), (b) and (c). Panel (a) shows the results obtained by using RF, and panel (b) is for a FCNN. A direct comparison suggests that the FCNN performs much better than RF for this problem, as the hidden neurons are able to extract relevant features out of the original sample. We argue that, in particular, the FCNN is able to obtain the information most likely related to the moments of order two, which is crucial to classify correctly in the sub-Rayleigh regime.
@@
Finally, also for a fair comparison with SPADE, we helped the algorithm by performing a pre-computation of the second moments. Starting from the relative frequencies $f_{x^{\prime}y^{\prime}}^{\text{DI}}(N)$, the second moments can be estimated as follows:
$\langle x^{2}\rangle$ $\leftarrow\sum_{x^{\prime},y^{\prime}}f_{x^{\prime}y^{\prime}}^{\text{DI}}(N){x^{\prime}}^{2}\,,$ (30)
$\langle y^{2}\rangle$ $\leftarrow\sum_{x^{\prime},y^{\prime}}f_{x^{\prime}y^{\prime}}^{\text{DI}}(N){y^{\prime}}^{2}\,,$ (31)
@@
$\frac{1}{2}\langle(x+y)^{2}\rangle$ $\leftarrow\frac{1}{2}\sum_{x^{\prime},y^{\prime}}f_{x^{\prime}y^{\prime}}^{\text{DI}}(N)(x^{\prime}+y^{\prime})^{2}\,,$ (32)
$\frac{1}{2}\langle(x^{\prime}-y^{\prime})^{2}\rangle$ $\leftarrow\frac{1}{2}\sum_{x^{\prime},y^{\prime}}f_{x^{\prime}y^{\prime}}^{\text{DI}}(N)(x^{\prime}-y^{\prime})^{2}\,,$ (33)
for the diagonal one. In particular, we have used the estimates of the second moments in diagonal basis and used them as feature vectors in a RF. The results are displayed in the bottom panel (c) of Fig. 4, showing an increase in the accuracy compared to the FCNN fed with the relative frequencies. Moreover, it can be noticed that the fluctuations for each $\tilde{\sigma}$ are more contained compared to the previous cases: this is due to the fact that the second moments, which now are structured within the feature vector, are independent of $\tilde{\sigma}$ and hence they stabilise the computation.
In conclusion, DI supported by either RF and FCNN yields poor performances when compared against SPADE supported by RF. The accuracy of DI increases when the second moments are pre-computed from the DI relative frequencies and fed into a RF. This latter approach mimics the pre-processing implemented by SPADE, though in a computational rather than physical way. However, it also performs poorly in comparison with SPADE, especially for low photon number and severe blurring. This is due to the fact that computing the moments from the data obtained by DI yields a signal-to-noise ratio smaller than a direct measurement of SPADE Santamaria22 .
Figure 5: Confusion matrix of the output of the RF for multi-class classification, performed on SPADE vectors. In particular the figure refers to the case with an effective blurring factor $\tilde{\sigma}=10$ and with number of photon detection events $N=5000$. In this numerical experiment, half of the photons are measured in the diagonal basis [Eq. (29)] and the other half in the Cartesian basis [Eqs. (35)-(36)].